### v3

# Explain with Visual Keypoints Like a Real Mentor! A Benchmark for Multimodal Solution Explanation

Abstract
With the rapid advancement of mathematical reasoning capabilities in Large Language Models (LLMs), AI systems are increasingly being adopted in educational settings to support students’ comprehension of problem-solving processes. However, a critical component remains underexplored in current LLM-generated explanations: multimodal explanation. In real-world instructional contexts, human tutors routinely employ visual aids, such as diagrams, markings, and highlights, to enhance conceptual clarity. To bridge this gap, we introduce the multimodal solution explanation task, designed to evaluate whether models can identify visual keypoints, such as auxiliary lines, points, angles, and generate explanations that incorporate these key elements essential for understanding. To evaluate model performance on this task, we propose ME2, a multimodal benchmark consisting of 1,000 math problems annotated with visual keypoints and corresponding explanatory text that references those elements. Our empirical results show that, aside from recent large-scale open-source and closed-source models, most generalist open-source models, and even math-specialist models, struggle with the multimodal solution explanation task. This highlights a significant gap in current LLMs’ ability to reason and explain with visual grounding in educational contexts. We expect that the multimodal solution explanation task and the ME2 dataset will catalyze further research on LLMs in education and promote their use as effective, explanation-oriented AI tutors. Code and data will be released publicly.
Figure 1: A student solving a math problem often benefits from visual cues, such as lines, symbols, or highlights, that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions. To serve as effective educational assistants, machines must go beyond answer generation and emulate human-like explanation strategies by explicitly incorporating and referencing visual elements.
1 Introduction
The traditional one-to-many educational model (i.e., one teacher for multiple students) is gradually transitioning to one-to-one personalized tutoring systems and online learning (Mukul and Büyüközkan 2023). Recent developments in Multimodal Large Language Models (MLLMs) have opened new opportunities for effective learning, such as estimating question difficulty (Park et al. 2024), assisting teachers in curriculum planning (Hu et al. 2024), and supporting interactive tutoring systems (Chevalier et al. 2024). In particular, numerous studies (Liu et al. 2023; Uesato et al. 2022; Lu et al. 2023) have focused on enhancing the mathematical reasoning abilities of MLLMs. As a result, MLLMs have led many students to use them as tools when faced with mathematical questions (Pardos and Bhandari 2024).
Figure 2: An overview of the ME2 benchmark. The ME2 consists of multimodal problem–solution pairs curated from real-world educational settings, along with visual keypoints and explanation summaries generated through a Human–AI annotation.
Figure 3: We propose two subtasks to robustly analyze multimodal solution explanation capacity: (1) Visual Keypoint Identification, which challenges machines to recognize visual keypoints useful for subsequent explanation, and (2) Keypoint-based Explanation Generation, which requires models to generate explanations that explicitly reference the identified visual keypoints.
However, from a student’s perspective, relying solely on the reasoning footprints of MLLMs may not always be the best way to understand problems (Pardos and Bhandari 2024; Jia et al. 2024). One might wonder what distinguishes a broadly comprehensible explanation from a solution that merely yields the correct answer, for either a human or a model? A critical factor is the use of visual cues.
In actual educational settings, Dual Coding Theory (DCT) naturally occurs, providing effective learning opportunities for students (Paivio 2013, 1990). According to DCT, combining verbal and visual information enhances student comprehension (Clark and Paivio 1991). As illustrated in Figure 1, human mentors often use visual scaffolding, such as annotated diagrams or highlighted keypoints on a blackboard, to foster intuitive understanding (Arcavi 2003; Stylianou 2010; Lee, Park, and Park 2024). In contrast, current AI models lack the capacity to generate such visual explanations. Moreover, existing datasets for training and evaluation are insufficient for developing models capable of producing these forms of multimodal instructional support (Hendrycks et al. 2021; Lu et al. 2023; Wang et al. 2024a).
To address these limitations, we introduce multimodal solution explanation, a novel task designed to evaluate and advance models’ ability to generate educationally effective, visually grounded mathematical explanations. To benchmark performance on multimodal solution explanation, we propose Multimodal Explanations for Mathematics Education (ME2) benchmark. The multimodal solution explanation task requires (1) identifying visual keypoints that are not present in the original problem but crucial for understanding (e.g., lines, angles, annotations) and (2) generating explanatory text that explicitly refers to them. The ME2 includes not only the original problem images and text, but also annotations of the visual keypoints that illustrate how the solution images differ from the originals, as well as keypoint-based solution explanations. Notably, we emphasize the educational value of ME2, as it is derived from real-world instructional contexts.
Experiments on ME2 demonstrate that, while closed-source models show potential on multimodal solution explanation, most current open-source generalist models still struggle to reliably identify visual keypoints. Even mathematical models, which are fine-tuned on mathematical domains, remain primarily focused on problem-solving and often do not produce sufficiently structured outputs to highlight critical visual details. These findings underscore a critical opportunity to steer future model development toward more effective visual grounding in educational contexts. We believe that ME2 will catalyze research aimed at strengthening visual grounding for mathematical reasoning and advancing multimodal solution explanation models capable of serving as effective, student-friendly educational mentors.
Our contributions are as follows:
  1. 1.
    A multimodal solution explanation task that supports students’ educational comprehension by identifying critical visual keypoints and generating explanatory text that explicitly references them.
  2. 2.
    A ME2 benchmark, rooted in authentic educational contexts, to rigorously assess multimodal solution explanation performance and facilitate further research on model-based explanations in real-world settings.
  3. 3.
    Extensive experimental evaluations of state-of-the-art MLLMs on multimodal solution explanation task, highlighting current limitations in recognizing and leveraging crucial visual keypoints to support effective learning.
2 ME2 Benchmark
ME2 is a multimodal solution explanation benchmark consisting of 1,000 instances. Each of which contains a problem text ($T_{p}$), a problem image ($I_{p}$), an explanatory solution text ($T_{s}$), a solution image ($I_{s}$), and visual keypoints ($VK$) that highlight how the solution image differs from the original, along with a concise summary of the explanation ($T_{s}^{tldr}$) to fix the solution direction and assess alignment between keypoints and their corresponding explanations, even when multiple valid solution exist.
To create a benchmark that can assess the multimodal solution explanation capabilities of MLLMs, we define the visual keypoint $VK$ and summary of the explanation $T_{s}^{tldr}$ in Section 2.1. An overview of ME2 and its construction is illustrated in Figure 2.
2.1 Benchmark Construction
In-house Data Curation.
We extract 1,000 instances of multimodal problem–solution pairs $\langle T_{p},\,I_{p},\,T_{s},\,I_{s}\rangle$ from an in-house mathematics education platform. All instances were authored by domain experts in mathematics to support effective student learning and are derived from materials authentically used in real-world educational contexts. The instances are written in Korean and span middle- to high-school levels, with a primary focus on geometry and graph theory. All benchmark data were carefully curated to ensure compliance with copyright regulations.
Total problem–solution pairs 1,000
- Geometry 763
   - Multiple-choice questions 464
   - Short-answer questions 299
- Geometry 237
   - Multiple-choice questions 141
   - Short-answer questions 96
Average number of $VK$ 3.8
Maximum words in $T_{p}$ 211
Maximum words in $T_{s}$ 361
Maximum words in $vk_{n}$ 45
Maximum words in $T_{s}^{tldr}$ 123
Average words in $T_{p}$ 53.1
Average words in $T_{s}$ 102.4
Average words in $vk_{n}$ 12.2
Maximum words in $T_{s}^{tldr}$ 35.7
Table 1: Statistics of the ME2 benchmark, including problem subjects and types, visual keypoints, and word counts.
To benchmark the models’ ability to recognize visual keypoints, we ensure that each solution image $I_{s}$ is derived from the corresponding problem image $I_{p}$ by adding only new elements such as points, angles, lines, regions, and symbols while preserving the original structure. For instance, in Figure 2, points are added to problem image $I_{p}$ to produce solution image $I_{s}$. This setup allows us to accurately evaluate whether a model can identify the critical visual keypoints and effectively incorporate them into its explanation.
We strictly curate the dataset to single-image math problems with either multiple-choice or short-answer formats. We focus on the domains of geometry and graph to ensure that visual context is essential for solving each problem. Each sample in ME2 consists of two natural language texts (the problem text $T_{p}$ and the solution text $T_{s}$) and two RGB images (the problem image $I_{p}$ and the solution image $I_{s}$).
Annotation Process.
To create the textual visual keypoints, we streamline the simple yet labor-intensive task of comparing problem and solution images by using GPT-4o (Achiam et al. 2023) as an auxiliary tool. The model produces an initial set of keypoints $\{vk^{ai}_{1},\dots,vk^{ai}_{n}\}\in VK^{ai}$, which four annotators, each holding a bachelor’s degree in science or engineering, verify and refine for accuracy and compliance with our formatting guidelines:
Any element that is newly added or modified, including points, lines, angles, regions, or symbols such as parallel marks, congruence marks, right-angle marks, or length labels, must be recorded in the format {element: description}, where element identifies the visual feature and description explains how it is introduced with reference to surrounding features.
Once the verified keypoints are fixed, we generate a brief, keypoint-aligned summary of each solution text, $T_{s}^{tldr}$. This summary captures the essential reasoning trajectory so that, even when a problem permits multiple valid approaches, a single, unambiguous set of visual keypoints remains authoritative. As with the keypoints, AI tool produces the initial draft, and human annotators review and refine it for clarity and consistency. Consequently, each ME2 instance is represented as $\langle T_{p},I_{p},T_{s},VK,T_{s}^{\text{tldr}}\rangle$. Further details are provided in the Supplementary Materials.
Figure 4: Topic coverage of geometry and graph across 17 chapters in the ME2 benchmark.
2.2 Data Analysis
The ME2 benchmark consists of 1,000 problem–solution pairs: 763 (76.3%) geometry problems and 237 (23.7%) graph problems. Among these, 605 (60.5%) are multiple-choice questions and 395 (39.5) are short-answer questions. It spans 17 chapters (see Figure 4) and 33 sections (see Supplementary Materials). On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations. These keypoints fall into four main categories: points, lines, regions, and symbols. The symbol category is further divided into parallel marks, equal-length marks, right-angle marks, and length-label marks. Additional statistical details about visual keypoints $VK$, and length statistics for the problem text $T_{p}$, the solution text $T_{s}$, and the visual keypoint components $vk_{n}$ are provided in Table 1.
3 Task Definition
We propose two tasks to evaluate a model’s multimodal solution explanation capability. Multimodal solution explanation requires identifying visual keypoints useful for subsequent explanation and generating explanations that explicitly reference them. These tasks offer a chance to assess how well the model understands multimodal educational elements and produces informative explanations along with visual keypoints. The two tasks are illustrated in Figure 3.
Visual Keypoint Identification.
The first task evaluates the model’s ability to identify visual keypoints that are crucial for comprehension. To prevent ambiguity arising from multiple solution paths to each problem, we supply the model with a summary of solution text $T_{s}^{tldr}$, ensuring it focuses on a single solution path and enabling precise evaluation of its ability to identify the corresponding visual keypoints. Given a problem image $I_{p}$, its text $T_{p}$, the correct answer, and the solution summary $T_{s}^{tldr}$, the model must select, from five candidate sets, the visual keypoints ($VK$) essential for understanding. The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem. We compute text similarity using Qwen3 Embedding (Zhang et al. 2025).
Keypoint-based Explanation Generation.
The second task evaluates whether the model can effectively generate explanatory text grounded in the appropriate visual keypoints. We provide the set of visual keypoints ($VK$) to eliminate ambiguity and guide the model toward a single correct reasoning path. Given a problem consisting of an image $I_{p}$, text $T_{p}$, and problem’s answer, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that refers to the relevant visual elements.
Problem-Solving (Acc)
Model Params Geometric Graph Overall
Molmo 7B 0.248 0.194 0.235
LLaVA-1.6 7B 0.147 0.127 0.142
Qwen2-VL 7B 0.274 0.215 0.260
Qwen2.5-VL 7B 0.316 0.224 0.294
Qwen2.5-VL 72B 0.430 0.300 0.399
Math-PUMA 7B 0.258 0.194 0.243
URSA 8B 0.055 0.068 0.058
Math-LLaVA 13B 0.202 0.152 0.190
GPT-4o - 0.274 0.211 0.259
Gemini 2.0 F - 0.481 0.291 0.436
Table 2: Experimental results on the Solution Recognition toy task from ME2. The best scores are shown in bold, and the second-best scores are underlined.
4 Experiments
Models.
We evaluate three categories of MLLMs: Open-source Generalist (OG), Open-source Mathematical specialist (OM), and Closed-source Generalist (CG). The OG models includes widely used models such as Molmo 7B (Deitke et al. 2024), LLaVA-1.6 7B (Liu et al. 2024), Qwen 2-VL (Wang et al. 2024b), and Qwen 2.5-VL in both 7B and 72B versions (Bai et al. 2025). The CG models comprise GPT-4o (Achiam et al. 2023) and Gemini 2.0 Flash (Google DeepMind 2024). Finally, the OM models, specialized for mathematics, include Math-PUMA 7B (Zhuang et al. 2024), URSA 8B (Luo et al. 2025), and Math-LLaVA 13B (Shi et al. 2024). Details of the experimental setup and prompts are provided in the Supplementary Materials.
4.1 Toy: Solution Recognition
Inspired by real-world tutoring settings where tutors typically know the solution path, the multimodal solution explanation tasks are designed under the assumption that the model has access to the correct solution, either in the form of a summary solution or a set of visual keypoints. To assess whether the model understands how to solve problems, we first conduct a preliminary study on ME2 before proceeding with the multimodal solution explanation task.
Metrics.
ME2 consists of both multiple-choice and short-answer problems. We report accuracy following the MathVista evaluation protocol (Lu et al. 2023).
Results.
Table 2 shows the accuracy of each model on the Solution Recognition toy task. The 7B OG baseline struggles, while the 72B model performs second best. Somewhat unexpectedly, the OM models perform worse than the OG models, likely due to aggressive fine-tuning that hinders instruction-following capabilities. Math-PUMA, despite being fine-tuned for mathematics on Qwen2-VL, also underperforms compared to the original models. Among the CG baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall. These results indicate that most open-source models struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task
4.2 Visual Keypoint Identification
Metrics.
We evaluate baselines’ VK identification performance using accuracy under a multiple-choice setting.
Results.
Table 3 summarizes performance on the VK Identification task. Among the OG baselines, the 7B models still face considerable difficulty, while the 72B model remains the second-best performer, similar to the Solution Recognition toy task. In contrast, the OM models perform at or below chance level (Acc = 0.20), even worse than on the toy task, suggesting a fundamental inability to detect relevant visual cues or a loss of instruction-following capability. Among the CG models, GPT-4o shows the most noticeable improvement despite its earlier underperformance on the toy task, while Gemini consistently achieves the highest performance. These findings indicate that although most open-source models still struggle despite having access to the solution, CG models and larger OG models can reasonably identify the visual keypoints.
VK Identification (Acc)
Model Params Geometric Graph Overall
Molmo 7B 0.253 0.312 0.267
LLaVA-1.6 7B 0.260 0.283 0.265
Qwen2-VL 7B 0.273 0.371 0.296
Qwen2.5-VL 7B 0.363 0.532 0.403
Qwen2.5-VL 72B 0.486 0.696 0.536
Math-PUMA 7B 0.194 0.219 0.200
URSA 8B 0.028 0.034 0.029
Math-LLaVA 13B 0.218 0.215 0.217
GPT-4o - 0.418 0.646 0.472
Gemini 2.0 F - 0.529 0.726 0.576
Table 3: Experimental results for the Visual Keypoint Identification task on ME2, where models are evaluated on their ability to select the correct keypoints from multiple-choices.
Figure 5: Examples of reasoning processes and final predictions produced by Qwen2.5-VL 7B, Math-PUMA, and Gemini 2.0 Flash on the Visual Keypoint Identification task. Qwen2.5-VL demonstrates task understanding and reasoning but produces an incorrect answer, Math-PUMA lacks both, while Gemini 2.0 Flash demonstrates both and produces the correct answer.
4.3 Keypoint-based Explanation Generation
Metrics.
We evaluate the quality of the explanation using three criteria: (1) Correctness – whether the model’s reasoning is logically sound and leads to a valid solution; (2) Fidelity – whether the explanation aligns with the reasoning and intent of the reference, regardless of surface form; (3) Referencing – whether the explanation refers to the same key visual components (e.g. points, lines, etc) as the reference. Each criterion is rated on a 5-point Likert scale. We report results from both human evaluators (Zheng et al. 2023) and an LLM-based evaluator (Achiam et al. 2023). In addition, we report text similarity metrics, including BLEU, ROUGE, METEOR, and BERTScore (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2019), with further details provided in the Supplementary Materials.
Model Correctness Fidelity Referencing
Molmo 2.362 1.906 2.545
LLaVA-1.6 1.815 1.545 1.970
Qwen2-VL 1.844 1.623 1.898
[…]
Math-PUMA 1.770 1.609 1.581
URSA 1.148 1.140 1.155
Math-LLaVA 2.100 1.360 1.286
GPT-4o 3.784 3.153 3.892
Gemini 2.0 F 3.849 3.489 4.103
Table 4: LLM-based evaluation results for the Keypoint-based Explanation Generation task on ME2, rated on a 1-5 Likert scale across three criteria: (1) Correctness, assessing logical validity; (2) Fidelity, measuring alignment with the intent of the reference explanation; and (3) Referencing, evaluating the appropriate use of key visual elements.
Model Correctness Fidelity Referencing
[…]
Math-PUMA 1.041 1.041 1.037
Gemini 2.0 F 4.423 4.171 4.256
Table 5: Human evaluation results for the Keypoint-based Explanation Generation task, rated on a 1–5 Likert scale.
Results.
Table 4 presents the results of the explanation generation task. Although all models perform relatively well in terms of Correctness by aligning their explanations with the problem, many fail to follow the intended reasoning path (Fidelity) or to reference the given visual keypoints (Referencing). OG models generally struggle, whereas the Qwen2.5-VL series performs relatively well, showing improved performance with increasing model size. OM models still fail to generate coherent explanations or follow instructions. In contrast, CG models achieved the highest scores in this explanation task, indicating their strong capabilities in generating well-grounded explanations. As shown in Table 5, human evaluation results show a similar pattern, with strong correlations between human and LLM judgments as indicated by the Spearman coefficients (Zar 2005), which are 0.770 for Correctness, 0.783 for Fidelity, and 0.788 for Referencing, all statistically significant (p$<$0.05). These results suggest that while most open-source models still struggle, CG models and recent OG models like the Qwen2.5-VL series are capable of generating appropriate explanations.
5 Analyses
5.1 Qualitative Analysis
To analyze how the three categories of models differ in their outputs, we examined the results from representative models in each category: Qwen2.5-VL 7B (OG), Math-PUMA (OM), and Gemini 2.0 Flash (CG).
Visual Keypoint Identification
Figure 5 shows VK identification examples from three model categories. The Qwen2.5-VL tries to reason about which option provides the most informative visual keypoints for the task, yet its final selection is wrong. The Math-PUMA shows neither coherent reasoning nor a correct answer. In contrast, the Gemini 2.0 Flash quickly grasps the core of the instruction, analyzes candidate visual keypoints, and selects the option that best supports understanding the problem.
Figure 6: Examples of explanations from three models on the Keypoint-based Explanation Generation task, along with their evaluated scores in Correctness (C), Fidelity (F), and Referencing (R). Qwen2.5-VL starts with valid yet unaligned explanations, which soon become incorrect; Math-PUMA generates no explanation; Gemini 2.0 Flash generates solution-aligned explanations.
Keypoint-based Explanation Generation
Figure 6 presents explanation examples from three model categories. In this example, Qwen2.5-VL scored 2 for Correctness, 2 for Fidelity, and 3 for Referencing. Although the model is given keypoints, its approach is not aligned with the reference solution. Its explanation references visual cues such as angles, lengths, symbols, and triangles, which indicates partial alignment with the solution image. However, it ultimately misinterprets the problem and produces incorrect reasoning. Math-PUMA received a score of 1 across all metrics, as it only returned the final answer without any supporting explanation. In contrast, Gemini scored 5 in all metrics. It generated an explanation as if the keypoints were embedded in the solution image, following the same reasoning structure as the reference and correctly referring to the same key visual elements throughout the solution.
Model PS Only VKI Only PS $\cap$ VKI
[…]
Math-PUMA 19.7% 15.5% 4.6%
GPT-4o 14.6% 35.9% 11.3%
Gemini 2.0 F 20.5% 34.5% 23.1%
Table 6: Proportion (%) of cases where each model succeeds only on Problem-Solving (PS), only on Visual Keypoint Identification (VKI), or on both (PS $\cap$ VKI)
5.2 Problem Solving and VK Identification
Table 6 reports the proportion of cases where models succeed only on Problem-Solving (PS), only on Visual Keypoint Identification (VKI), or on both. Overall, generalist models tend to perform better on VKI than on PS alone, suggesting that most models are more capable of recognizing keypoints assuming the reasoning path is known, rather than solving the problems end-to-end. Notably, Qwen 72B and Gemini achieve over 20% accuracy in cases involving both PS and VKI, demonstrating relatively strong performance. However, since current models still struggle with problem-solving, improving multimodal mathematical reasoning is essential for advancing performance on both PS and VKI.
Figure 7: Attention maps from various layers of open-source MLLMs on the Visual Keypoint Identification task. While the models attend visually to the problem image, they fail to focus on the keypoints that are most relevant for explanation.
5.3 Do MLLMs Attend to Visual Keypoints?
To analyze whether current open-source models can accurately identify visual keypoints in images, we examined the attention maps of OG and OM models on the multimodal solution explanation tasks. Figure 7 illustrates the attention patterns of the OG and OM models. Both models show a tendency to attend to both global and local regions of the input image. However, despite the explicit inclusion of visual keypoints in the options and summary solution of the Visual Keypoint Identification task, neither model effectively focuses on the relevant visual regions. A similar pattern is observed in the Keypoint-based Explanation Generation task. These findings suggest that current open-source models still struggle to visually attend to the keypoints, even when they are essential for understanding or explaining the problem.
6 Related Works
Language Models for Education.
Recent advances in Large Language Models (LLMs) (Brown et al. 2020) have sparked significant interest in educational applications, particularly in personalized problem recommendation (Park et al. 2024), automated tutoring (Chevalier et al. 2024), and the provision of tailored feedback and customized curricula (Hu et al. 2024; Macina et al. 2023; Feng, Wang, and Sun 2023). For effective education, research suggests that combining textual and visual information enhances comprehension and memory more effectively than using text alone (Arcavi 2003; Stylianou 2010; Lee, Park, and Park 2024). As Clark and Paivio (1991) explains, dual representations supply multiple retrieval cues and cultivate richer mental models. Building on this insight, we introduce the multimodal solution explanation task to enable LLMs to offer learners more comprehensive learning opportunities. This task enables the model to pinpoint the visual keypoints essential for students’ comprehension and to generate explanations grounded in those keypoints, thereby delivering more comprehensive educational support and ultimately improving the overall quality of their learning experience.
Multimodal Mathematical Benchmarks.
Current LLMs show strong performance on mathematical problems, making them valuable tools for students (Zhuang et al. 2024; Luo et al. 2025). To evaluate these models, traditional mathematical benchmarks (Cobbe et al. 2021; Hendrycks et al. 2021) have been crucial in assessing reasoning capabilities. With the growing multimodal capabilities of LLMs, benchmarks such as MathVista (Lu et al. 2023), Math-Vision (Wang et al. 2024a), and MathVerse (Zhang et al. 2024) have extended this evaluation to image-based math problems. Recent efforts like OlympiadBench (He et al. 2024) and MM-MATH (Sun et al. 2024) further assess models not just on final answers but also on their reasoning processes. However, most existing benchmarks focus solely on problem-solving, overlooking educational objectives. To address this gap, we introduce ME2, a new benchmark designed to evaluate a model’s multimodal solution explanation capability. Together, ME2 and the multimodal solution explanation task are expected to advance educational research and foster the development of more effective teaching models.
7 Disscusion and Conclusion
We introduce the multimodal solution explanation task, which requires core mathematics-teaching capabilities for generating intuitive and insightful explanations. To benchmark these capabilities, we propose ME2 benchmark, which consists of two subtasks: (1) Visual Keypoint Identification, where models must select visual keypoints that are essential for understanding and explaining the problem, and (2) Keypoint-based Explanation Generation, which requires models to produce explanations grounded in those visual keypoints that support intuitive reasoning.
Our experimental results show that current MLLMs struggle even with the problem-solving aspects of the ME2 tasks. Furthermore, even under the assumption that the model already knows the solution path, similar to a human tutor, we found that most open-source models still fail to accurately identify visual keypoints, with the exception of closed-source models and recent open-source models. Surprisingly, even math-specialized models perform worse in this aspect. The performance gap becomes even more pronounced in the explanation generation task. To perform well on the multimodal solution explanation task, future open-source models may need the following advancements:
  1. 1.
    Enhanced visual grounding mechanisms that link visual elements to multimodal mathematical reasoning.
  2. 2.
    Advanced solution-path tracking capabilities for multimodal explanations.
We believe this work will inspire future research along two key paths: (1) enhancing models’ mathematical reasoning capabilities through improved visual grounding, and (2) leveraging visual grounding in educational contexts to develop AI systems that can serve as real-world mentors, ultimately providing clear, multimodal, and student-friendly explanations. The code and data will be released publicly.
A Benchmark for Multimodal Solution Explanation
Supplementary Material
Appendix A Overview
This supplementary material includes the following sections: (1) Detailed experimental settings used in our benchmark, (2) Details of our benchmark, and (3) Prompts used for model response generation
Appendix B Experiment Details
In this study, experiments were conducted using the LMMs-eval repository (Li et al. 2024). This repository provides a comprehensive framework for evaluating multi-modal models across various tasks.
B.1 Evaluated Models
In all LLM-based evaluations, we used the gpt-4o-2024-08-06 endpoint. When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^1] https://huggingface.co/models. The following model parameters were used for each model:
  • Math-PUMA: Math-PUMA/Math-PUMA_Qwen2VL-7B
  • URSA: URSA-MATH/URSA-RM-8B
  • Math-LLaVA: Zhiqiang007/Math-LLaVA
  • llava-1.6: llava-hf/llava-v1.6-mistral-7b-hf
  • Qwen2-VL-7B: Qwen/Qwen2-VL-7B-Instruct
  • Qwen2.5-VL-7B: Qwen/Qwen2.5-VL-7B-Instruct
  • Qwen2.5-VL-72B: Qwen/Qwen2.5-VL-72B-Instruct
  • Molmo: allenai/Molmo-7B-D-0924
These models were evaluated in our benchmark, which included tasks designed to assess both visual understanding and textual explanation capabilities. The selection of models spans a range of architectures and performance levels, providing insights into current advancements in multi-modal learning.
B.2 Computational Resources
For closed-source models, we used the OpenAI API and Gemini Developer API to infer the output of GPT-4o and Gemini 2.0. For open-source models such as Math-PUMA, URSA, MathLLaVA, LLaVA, QWEN-2-VL, QWEN-2.5-VL, and Molmo, inference was performed using a NVIDIA A6000 48GB GPU. While the exact inference speed varies depending on the task, model, and lengths of prompts and responses, a query takes about 50 seconds to be answered.
B.3 Human Evaluation Setup
Three evaluators, all holding a bachelor’s or master’s degree in engineering, are assessing AI model outputs for 80 problems. Specifically, they are evaluating the results produced by the Math-PUMA, Qwen2.5-VL, and Gemini 2.0 Flash models, which represent the OM, OG, and CG categories. Each criterion is being rated on a five-point Likert scale.
B.4 Evaluation on automatic metrics
We report additional evaluation results for Keypoint-based Explanation Generation task using automatic metrics, including BLEU, ROUGE, METEOR, and BERTScore (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2019). The detaileed scores can be found in Table A.
Appendix C Benchmark Details
Our ME2 benchmark consists of a total of 17 chapters and 33 sections as shown in Table B.
Appendix D Prompts
This section compiles all the prompts used in our experiments. The prompts shown in Figure A, Figure B, and Figure C are used to generate model outputs for the Solution Recognition toy task, Visual Keypoint Identification, and Keypoint-based Explanation Generation tasks, respectively. For the Keypoint-based Explanation Generation task, model responses are evaluated using the prompts in Figure D.
Appendix E Ethical Considerations
In this paper, we introduce the ME2 benchmark, curated from an in-house mathematics education platform. All materials were reviewed to ensure compliance with copyright regulations. Data annotation was performed by bilingual annotators with undergraduate degrees, ensuring sufficient mathematical problem-solving skills.
Model Params BLEU-2 BLEU-4 ROUGE-L METEOR BertScore
OG Molmo 7B 0.158 0.067 0.187 0.310 0.842
LLaVA-1.6 7B 0.130 0.059 0.176 0.287 0.835
Qwen2-VL 7B 0.176 0.087 0.237 0.288 0.854
Qwen2.5-VL 7B 0.099 0.038 0.193 0.284 0.819
Qwen2.5-VL 72B 0.097 0.043 0.190 0.316 0.821
OM Math-PUMA 7B 0.006 0.002 0.119 0.058 0.818
URSA 8B 0.020 0.006 0.079 0.075 0.735
Math-LLaVA 13B 0.112 0.057 0.147 0.252 0.817
CG Gemini 2.0 Flash - 0.149 0.083 0.210 0.367 0.842
GPT-4o - 0.095 0.045 0.161 0.301 0.815
Table A: Experimental results of automated evaluation metrics for the Keypoint-based Explanation Generation task on ME2.
Chapter Title Section Title
Basics of Geometry Basic Geometric Figures
Construction and Congruence
Coordinate Plane and Graphs Coordinate Plane and Graphs
Differential Calculus Differentiation of Various Functions
Differentiation Derivative and Derivative Function
Applications of Derivatives
Equations and Inequalities Quadratic Equations and Functions
Equations of Geometric Figures Transformations of Figures
Equation of a Circle
Coordinate Plane
Equations of Straight Lines
Exponential and Logarithmic Functions Exponential and Logarithmic Functions
Functions Linear Functions and Their Graphs
Functions
Rational and Irrational Functions
Relationship Between Linear Functions and Equations
Integral Calculus Applications of Definite Integrals
Various Integration Techniques
Integration Indefinite and Definite Integrals
Applications of Definite Integrals
Plane and Solid Figures Properties of Solid Figures
Properties of Plane Figures
Properties of Circles Circle and Line
Inscribed Angles
Properties of Figures Properties of Quadrilaterals
Properties of Triangles
Quadratic Functions Graph of the Quadratic Function $y=ax^{2}+bx+c$
Similarity and the Pythagorean Theorem Pythagorean Theorem
Similarity of Figures
Applications of Similarity
Trigonometric Functions Meaning and Graphs of Trigonometric Functions
Law of Sines and Law of Cosines
Trigonometric Ratios Trigonometric Ratios
Applications of Trigonometric Ratios
Table B: Overview of the 17 chapters and their corresponding 33 sections covered in the dataset.
You should choose a set of visual elements from the multiple-choice options (A, B, C, D, or E) that best reflect how a teacher would visually guide a student to understand and solve the problem. Problem: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b. Answer: 19 The solution process for the problem is as follows: Count the possible straight lines formed by selecting pairs of points, then count the rays formed by considering directionality. Add both counts to find the total. A.- line AE: A line connecting point A and E B.- Symbol a: Represents the line extending from the upper left to the lower right
- Symbol b: Represents the line extending from the lower left to the upper right
- Symbol c: Represents the horizontal line
C.- Line AB: A line extended from side AB of the hexagon
- Line BC: A line extended from side BC of the hexagon
- Line CD: A line extended from side CD of the hexagon
- Line DE: A line extended from side DE of the hexagon
- Line EF: A line extended from side EF of the hexagon
- Line FA: A line extended from side FA of the hexagon
D.- Auxiliary line BD: A line segment connecting point B and point D is added E.- Line AE: A straight line connecting points A and E
- Line BE: A straight line connecting points B and E
- Line CE: A straight line connecting points C and E
- Line DE: A straight line connecting points D and E
Based on this reasoning guidance, select **only one** of the option (A, B, C, D, or E) whose visual elements would be most helpful for students in understanding the problem and its solution.
Think carefully about how the selected visual elements support the reasoning process. You may briefly explain your thinking, but your response **must end** with the following format: The final answer is: A, B, C, D, or E IMPORTANT!! Your final response must END with the format.
Figure A: Prompt used for the Visual Keypoint Identification task in the ME2 benchmark. Prompt inputs are boldfaced.
Q: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b. ### Answer: 19 ### ### Difference between the original image and the solution image ### Line AE: A straight line connecting points A and E
Line BE: A straight line connecting points B and E
Line CE: A straight line connecting points C and E
Line DE: A straight line connecting points D and E
You are a math teacher helping students understand how to solve problems clearly and effectively. Given a problem description, problem image and a list of key elements introduced or highlighted in the solution image, write an educational explanation that helps students. Additionally, this problem is a problem of Functions/Linear Functions and Their Graphs chapter. You should explain the problem in the context of the chapter and section. Make sure to reference both the original components from the problem image and any new annotations, highlights, or added elements from the solution image to enhance understanding. ### OUTPUT Example: {
solution_text:
Figure B: Prompt used for the Keypoint-based Explanation Generation task in the ME2 benchmark. It is designed to generate educationally effective explanations for the given math problem. Prompt inputs are boldfaced.
You are a math solver. For the problem below, **your task is ONLY to output the final answer** in one line. **Do NOT provide any explanation, steps, or clarification. Just write the answer.** Problem: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b. Again, only return the final answer. Any additinal text will be considered incorrect.
Figure C: Prompt used for the Solution Recognition toy task in the ME2 benchmark. It is designed to generate an answer to the given math problem. Prompt inputs are boldfaced.
You are evaluating the quality of an AI-generated explanation for a math problem involving geometry or graph-based reasoning.
You will be given two texts:
1. A reference explanation written by a human teacher.
2. An AI-generated explanation written by a model.
Your task is to compare the two explanations and assess how accurately and effectively the AI-generated explanation captures the key geometric concepts and reasoning presented in the reference.
Please evaluate the model’s explanation and provide four scores based on the criteria below:
### Scoring Criteria
1. Correctness
- Does the reasoning presented by the model make sense and help solve the problem appropriately?
- Rate on a Likert scale: **1, 2, 3, 4, or 5**
2. Reference Alignment
- Does the model follow the same logical reasoning and intent as the reference explanation, even if the wording differs?
- Rate on a Likert scale: **1, 2, 3, 4, or 5**
3. Use of Key Visual Elements
- Does the AI explanation refer to the same critical visual components (e.g., points, lines, angles, shapes) as the reference?
- Alternative terminology is acceptable if it clearly refers to the same element or serves the same purpose.
- Rate on a Likert scale: **1, 2, 3, 4, or 5**
4. Mathematical Accuracy
- Does the model produce the same final answer as the reference?
- If the model’s answer is correct, give a score of **1**.
- If the model’s answer is incorrect, give a score of **0**.
### Output Format
Important: Report your rating using the exact format below:
Rating: [[x, y, z, w]]
— where ‘x‘ is your score for correctness, ‘y‘ for reference alignment, ‘z‘ for use of visual elements, and ‘w‘ for mathematical accuracy.
Figure D: GPT evaluation prompt used to assess model outputs for the Keypoint-based Explanation Generation task in ME2. Prompt inputs are boldfaced.

### v4

@@
 To evaluate model performance on this task, we propose ME2, a multimodal benchmark consisting of 1,000 math problems annotated with visual keypoints and corresponding explanatory text that references those elements.
-Our empirical results show that, aside from recent large-scale open-source and closed-source models, most generalist open-source models, and even math-specialist models, struggle with the multimodal solution explanation task.
-This highlights a significant gap in current LLMs’ ability to reason and explain with visual grounding in educational contexts.
+Our empirical results show that current models struggle to identify visual keypoints.
+In the task of generating keypoint-based explanations, open-source models also face notable difficulties.
+This highlights a significant gap in current LLMs’ ability to perform mathematical visual grounding, engage in visually grounded reasoning, and provide explanations in educational contexts.
 We expect that the multimodal solution explanation task and the ME2 dataset will catalyze further research on LLMs in education and promote their use as effective, explanation-oriented AI tutors.
-Code and data will be released publicly.
+Archivehttps://me2-benchmark.github.io
 Figure 1: A student solving a math problem often benefits from visual cues, such as lines, symbols, or highlights, that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions.
@@
 1 Introduction
-The traditional one-to-many educational model (i.e., one teacher for multiple students) is gradually transitioning to one-to-one personalized tutoring systems and online learning (Mukul and Büyüközkan 2023).
-Recent developments in Multimodal Large Language Models (MLLMs) have opened new opportunities for effective learning, such as estimating question difficulty (Park et al.
-2024), assisting teachers in curriculum planning (Hu et al.
-2024), and supporting interactive tutoring systems (Chevalier et al.
-2024).
-In particular, numerous studies (Liu et al.
-2023; Uesato et al.
-2022; Lu et al.
-2023) have focused on enhancing the mathematical reasoning abilities of MLLMs.
-As a result, MLLMs have led many students to use them as tools when faced with mathematical questions (Pardos and Bhandari 2024).
-Figure 2: An overview of the ME2 benchmark.
-The ME2 consists of multimodal problem–solution pairs curated from real-world educational settings, along with visual keypoints and explanation summaries generated through a Human–AI annotation.
-Figure 3: We propose two subtasks to robustly analyze multimodal solution explanation capacity: (1) Visual Keypoint Identification, which challenges machines to recognize visual keypoints useful for subsequent explanation, and (2) Keypoint-based Explanation Generation, which requires models to generate explanations that explicitly reference the identified visual keypoints.
-However, from a student’s perspective, relying solely on the reasoning footprints of MLLMs may not always be the best way to understand problems (Pardos and Bhandari 2024; Jia et al.
-2024).
+The traditional one-to-many educational model (i.e., one teacher for multiple students) is gradually transitioning to one-to-one personalized tutoring systems and online learning (mukul2023digital).
+Recent developments in Multimodal Large Language Models (MLLMs) have opened new opportunities for effective learning, such as estimating question difficulty (park2024large), assisting teachers in curriculum planning (hu2024teaching), and supporting interactive tutoring systems (chevalier2024language).
+In particular, numerous studies (liu2023mathematical; uesato2022solving; lu2023mathvista) have focused on enhancing the mathematical reasoning abilities of MLLMs.
+As a result, MLLMs have led many students to use them as tools when faced with mathematical questions (pardos2024chatgpt).
+However, from a student’s perspective, relying solely on the reasoning footprints of MLLMs may not always be the best way to understand problems (pardos2024chatgpt; jia2024comparison).
 One might wonder what distinguishes a broadly comprehensible explanation from a solution that merely yields the correct answer, for either a human or a model?
 A critical factor is the use of visual cues.
-In actual educational settings, Dual Coding Theory (DCT) naturally occurs, providing effective learning opportunities for students (Paivio 2013, 1990).
-According to DCT, combining verbal and visual information enhances student comprehension (Clark and Paivio 1991).
-As illustrated in Figure 1, human mentors often use visual scaffolding, such as annotated diagrams or highlighted keypoints on a blackboard, to foster intuitive understanding (Arcavi 2003; Stylianou 2010; Lee, Park, and Park 2024).
+In actual educational settings, Dual Coding Theory (DCT) naturally occurs, providing effective learning opportunities for students (paivio2013imagery; paivio1990mental).
+According to DCT, combining verbal and visual information enhances student comprehension (clark1991dual).
+As illustrated in Figure 1, human mentors often use visual scaffolding, such as annotated diagrams or highlighted keypoints on a blackboard, to foster intuitive understanding (arcavi2003role; stylianou2010teachers; lee2024vistavisualintegratedtailored).
 In contrast, current AI models lack the capacity to generate such visual explanations.
-Moreover, existing datasets for training and evaluation are insufficient for developing models capable of producing these forms of multimodal instructional support (Hendrycks et al.
-2021; Lu et al.
-2023; Wang et al.
-2024a).
-To address these limitations, we introduce multimodal solution explanation, a novel task designed to evaluate and advance models’ ability to generate educationally effective, visually grounded mathematical explanations.
+Moreover, existing datasets (hendrycks2021measuring; lu2023mathvista; wang2024measuring) focus solely on problem-solving and overlook educational objectives, making them insufficient for developing models capable of providing such forms of multimodal instructional support.
+To address these limitations, we introduce multimodal solution explanation, a novel task that aims to enhance models’ capacity to generate educationally effective and visually grounded mathematical explanations.
+In this task, models are required to (1) identify visual keypoints that are not present in the original problem but are crucial for understanding (e.g., lines, angles, annotations), and (2) generate explanatory text that explicitly refers to them.
 To benchmark performance on multimodal solution explanation, we propose Multimodal Explanations for Mathematics Education (ME2) benchmark.
-The multimodal solution explanation task requires (1) identifying visual keypoints that are not present in the original problem but crucial for understanding (e.g., lines, angles, annotations) and (2) generating explanatory text that explicitly refers to them.
-The ME2 includes not only the original problem images and text, but also annotations of the visual keypoints that illustrate how the solution images differ from the originals, as well as keypoint-based solution explanations.
-Notably, we emphasize the educational value of ME2, as it is derived from real-world instructional contexts.
-Experiments on ME2 demonstrate that, while closed-source models show potential on multimodal solution explanation, most current open-source generalist models still struggle to reliably identify visual keypoints.
-Even mathematical models, which are fine-tuned on mathematical domains, remain primarily focused on problem-solving and often do not produce sufficiently structured outputs to highlight critical visual details.
-These findings underscore a critical opportunity to steer future model development toward more effective visual grounding in educational contexts.
-We believe that ME2 will catalyze research aimed at strengthening visual grounding for mathematical reasoning and advancing multimodal solution explanation models capable of serving as effective, student-friendly educational mentors.
+The ME2 includes not only the problem and its solution, but also annotations of the visual keypoints that serve as visual cues necessary to explain the solution, as well as keypoint-based explanatory texts aligned with them.
+Notably, we emphasize that ME2 goes beyond simple problem-solving, offering educational value in addressing the previously unexplored dimension of multimodal solution explanation.
+Experiments on ME2 demonstrate that current MLLMs struggle to reliably identify visual keypoints.
+While closed-source models show potential in generating explanations grounded in visual keypoints, open-source generalist and math-specialized models show limited ability in this aspect.
+This suggests that current models largely fail to achieve robust mathematical visual grounding and visually grounded reasoning in educational contexts.
+We believe that ME2 will catalyze research toward strengthening mathematical visual grounding and reasoning, and advancing models that can serve as effective and student-friendly educational mentors.
 Our contributions are as follows:
@@
     Extensive experimental evaluations of state-of-the-art MLLMs on multimodal solution explanation task, highlighting current limitations in recognizing and leveraging crucial visual keypoints to support effective learning.
+Figure 2: An overview of the ME2 benchmark.
+The ME2 consists of multimodal problem–solution pairs curated from real-world educational settings, along with visual keypoints and explanation summaries generated through a Human–AI annotation.
+Figure 3: We propose two subtasks to robustly analyze multimodal solution explanation capacity: (1) Visual Keypoint Identification, which challenges machines to recognize visual keypoints useful for subsequent explanation, and (2) Keypoint-based Explanation Generation, which requires models to generate explanations that explicitly reference the identified visual keypoints.
+2 Related Works
+Language Models for Education.
+Recent advances in Large Language Models (LLMs) (NEURIPS2020_1457c0d6) have sparked significant interest in educational applications, particularly in personalized problem recommendation (park2024large), automated tutoring (chevalier2024language), and the provision of tailored feedback and customized curricula (hu2024teaching; macina2023mathdial; feng2023citing).
+For effective education, research suggests that combining textual and visual information enhances comprehension and memory more effectively than using text alone (arcavi2003role; stylianou2010teachers; lee2024vistavisualintegratedtailored).
+As clark1991dual explains, dual representations supply multiple retrieval cues and cultivate richer mental models.
+Building on this insight, we introduce the multimodal solution explanation task to enable LLMs to offer learners more comprehensive learning opportunities.
+This task enables the model to pinpoint the visual keypoints essential for students’ comprehension and to generate explanations grounded in those keypoints, thereby delivering more comprehensive educational support and ultimately improving the overall quality of their learning experience.
+Mathematical Benchmarks.
+Current LLMs show strong performance on mathematical problems, making them valuable tools for students (zhuang2024math; luo2025ursa).
+To evaluate these models, traditional mathematical benchmarks (cobbe2021training; hendrycks2021measuring) have been crucial in assessing reasoning capabilities.
+With the growing multimodal capabilities of LLMs, benchmarks such as MathVista (lu2023mathvista), Math-Vision (wang2024measuring), and MathVerse (zhang2024mathverse) have extended this evaluation to image-based math problems.
+Recent efforts like OlympiadBench (he2024olympiadbench) and MM-MATH (sun2024mm) further assess models not just on final answers but also on their reasoning processes.
+However, most existing benchmarks focus solely on problem-solving, overlooking educational objectives.
+To address this gap, we introduce ME2, which advances beyond problem-solving to evaluate a model’s capacity to generate visually and logically coherent explanations and key visual cues that support effective instructional use.
 3 ME2 Benchmark
 ME2 is a multimodal solution explanation benchmark consisting of 1,000 instances.
-Each of which contains a problem text ($T_{p}$), a problem image ($I_{p}$), an explanatory solution text ($T_{s}$), a solution image ($I_{s}$), and visual keypoints ($VK$) that highlight how the solution image differs from the original, along with a concise summary of the explanation ($T_{s}^{tldr}$) to fix the solution direction and assess alignment between keypoints and their corresponding explanations, even when multiple valid solution exist.
+Each of which contains a problem text ($T_{p}$), a problem image ($I_{p}$), an explanatory solution text ($T_{s}$), a solution image ($I_{s}$), and visual keypoints ($VK$) that newly highlight elements crucial for understanding (e.g., lines, angles), and a concise explanation summary ($T_{s}^{tldr}$) to anchor the model’s explanatory solution direction.
 To create a benchmark that can assess the multimodal solution explanation capabilities of MLLMs, we define the visual keypoint $VK$ and summary of the explanation $T_{s}^{tldr}$ in Section 3.1.
@@
 All benchmark data were carefully curated to ensure compliance with copyright regulations.
+To benchmark the models’ ability to recognize visual keypoints, we ensure that each solution image $I_{s}$ is derived from the corresponding problem image $I_{p}$ by adding only new elements such as points, angles, lines, regions, and symbols while preserving the original structure.
+For instance, in Figure 2, points are added to problem image $I_{p}$ to produce solution image $I_{s}$.
+This setup allows us to accurately evaluate whether a model can identify the critical visual keypoints and effectively incorporate them into its explanation.
+We strictly curate the dataset to single-image math problems with either multiple-choice or short-answer formats.
+We focus on the domains of geometry and graph to ensure that visual context is essential for solving each problem.
+Each sample in ME2 consists of two natural language texts (the problem text $T_{p}$ and the solution text $T_{s}$) and two RGB images (the problem image $I_{p}$ and the solution image $I_{s}$).
+Annotation Process.
+To create the textual-form visual keypoints, we streamline the simple yet labor-intensive task of comparing problem and solution images by using GPT-4o (achiam2023gpt) as an auxiliary tool.
+The model produces an initial set of keypoints $\{vk^{\text{ai}}_{1},\dots,vk^{\text{ai}}_{n}\}\in VK^{\text{ai}}$, which four annotators, each holding a bachelor’s degree in science or engineering, verify and refine for precision and consistency with our annotation guidelines:
+Any element that is newly added or modified, including points, lines, angles, regions, or symbols such as parallel marks, congruence marks, right-angle marks, or length labels, must be recorded in the format {element: description}, where element identifies the visual feature and description explains how it is introduced with reference to surrounding features.
+Once the verified keypoints are fixed, human annotators and the AI tool jointly generate a brief, keypoint-aligned summary of each solution text ($T_{s}^{tldr}$).
+Since explanations may follow multiple valid paths (see Appendix), this summary anchors a single solution direction during model explanation generation, ensuring an unambiguous consensus set of visual keypoints.
+Finally, the entire benchmark was translated from Korean to English using an AI tool and then reviewed by two bilingual annotators.
+From a 10% subset, annotators achieved substantial agreement, with Cohen’s $\kappa$ of 0.84 (cohen1960coefficient), indicating strong reliability.
+Consequently, each ME2 instance is represented as $\langle T_{p},I_{p},T_{s},VK,T_{s}^{\text{tldr}}\rangle$.
 Total problem–solution pairs 1,000
@@
    - Short-answer questions 299
-- Geometry 237
+- Graph 237
    - Multiple-choice questions 141
    - Short-answer questions 96
-Average number of $VK$ 3.8
+Average number of VKVK 3.8
 Maximum words in $T_{p}$ 211
@@
 Maximum words in $T_{s}^{tldr}$ 35.7
-Table 1: Statistics of the ME2 benchmark, including problem subjects and types, visual keypoints, and word counts.
-To benchmark the models’ ability to recognize visual keypoints, we ensure that each solution image $I_{s}$ is derived from the corresponding problem image $I_{p}$ by adding only new elements such as points, angles, lines, regions, and symbols while preserving the original structure.
-For instance, in Figure 2, points are added to problem image $I_{p}$ to produce solution image $I_{s}$.
-This setup allows us to accurately evaluate whether a model can identify the critical visual keypoints and effectively incorporate them into its explanation.
-We strictly curate the dataset to single-image math problems with either multiple-choice or short-answer formats.
-We focus on the domains of geometry and graph to ensure that visual context is essential for solving each problem.
-Each sample in ME2 consists of two natural language texts (the problem text $T_{p}$ and the solution text $T_{s}$) and two RGB images (the problem image $I_{p}$ and the solution image $I_{s}$).
-Annotation Process.
-To create the textual visual keypoints, we streamline the simple yet labor-intensive task of comparing problem and solution images by using GPT-4o (Achiam et al.
-2023) as an auxiliary tool.
-The model produces an initial set of keypoints $\{vk^{ai}_{1},\dots,vk^{ai}_{n}\}\in VK^{ai}$, which four annotators, each holding a bachelor’s degree in science or engineering, verify and refine for accuracy and compliance with our formatting guidelines:
-Any element that is newly added or modified, including points, lines, angles, regions, or symbols such as parallel marks, congruence marks, right-angle marks, or length labels, must be recorded in the format {element: description}, where element identifies the visual feature and description explains how it is introduced with reference to surrounding features.
-Once the verified keypoints are fixed, we generate a brief, keypoint-aligned summary of each solution text, $T_{s}^{tldr}$.
-This summary captures the essential reasoning trajectory so that, even when a problem permits multiple valid approaches, a single, unambiguous set of visual keypoints remains authoritative.
-As with the keypoints, AI tool produces the initial draft, and human annotators review and refine it for clarity and consistency.
-Consequently, each ME2 instance is represented as $\langle T_{p},I_{p},T_{s},VK,T_{s}^{\text{tldr}}\rangle$.
-Further details are provided in the Supplementary Materials.
-Figure 4: Topic coverage of geometry and graph across 17 chapters in the ME2 benchmark.
+Table 1: Statistics of the ME2 benchmark, including problem subjects, types, and instance word counts.
 3.2 Data Analysis
 The ME2 benchmark consists of 1,000 problem–solution pairs: 763 (76.3%) geometry problems and 237 (23.7%) graph problems.
-Among these, 605 (60.5%) are multiple-choice questions and 395 (39.5) are short-answer questions.
-It spans 17 chapters (see Figure 4) and 33 sections (see Supplementary Materials).
+Among these, 605 (60.5%) are multiple-choice questions and 395 (39.5%) are short-answer questions.
+It spans 17 chapters (see Figure 4) and 33 sections (see Appendix).
 On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations.
@@
 4 Task Definition
-We propose two tasks to evaluate a model’s multimodal solution explanation capability.
-Multimodal solution explanation requires identifying visual keypoints useful for subsequent explanation and generating explanations that explicitly reference them.
-These tasks offer a chance to assess how well the model understands multimodal educational elements and produces informative explanations along with visual keypoints.
-The two tasks are illustrated in Figure 3.
+We propose two tasks to evaluate a model’s multimodal solution explanation capability, as illustrated in Figure 3.
+The first task requires (1) identifying visual keypoints useful for subsequent explanation and (2) generating explanations that explicitly reference them.
+For robust evaluation, we structured the tasks to isolate perceptual and reasoning subskills.
+Although the design abstracts away some real-world complexity, this two-stage setup still provides a clear and measurable step toward unified, open-ended reasoning.
 Visual Keypoint Identification.
 The first task evaluates the model’s ability to identify visual keypoints that are crucial for comprehension.
-To prevent ambiguity arising from multiple solution paths to each problem, we supply the model with a summary of solution text $T_{s}^{tldr}$, ensuring it focuses on a single solution path and enabling precise evaluation of its ability to identify the corresponding visual keypoints.
+Since a problem may have multiple valid solutions and the corresponding visual keypoints can vary, we provide the model with a solution summary ($T_{s}^{tldr}$) that anchors a single explanatory direction.
+To ensure that the evaluation focuses on keypoint identification rather than problem-solving, the correct answer is also provided.
+Additionally, because current models cannot reliably generate valid keypoints and open-ended scoring is ambiguous, we adopt a multiple-choice format for robust evaluation.
 Given a problem image $I_{p}$, its text $T_{p}$, the correct answer, and the solution summary $T_{s}^{tldr}$, the model must select, from five candidate sets, the visual keypoints ($VK$) essential for understanding.
 The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem.
-We compute text similarity using Qwen3 Embedding (Zhang et al.
-2025).
+We compute text similarity using Qwen3 Embedding (qwen3embedding).
 Keypoint-based Explanation Generation.
 The second task evaluates whether the model can effectively generate explanatory text grounded in the appropriate visual keypoints.
-We provide the set of visual keypoints ($VK$) to eliminate ambiguity and guide the model toward a single correct reasoning path.
+As in the first task, we provide visual keypoints ($VK$) to guide the model toward a single reasoning path and the correct answer to focus evaluation on keypoint-aligned explanation generation rather than problem-solving.
 Given a problem consisting of an image $I_{p}$, text $T_{p}$, and problem’s answer, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that refers to the relevant visual elements.
+Figure 4: Topic coverage of geometry and graph across 17 chapters in the ME2 benchmark.
+5 Experiments
+Models.
+We evaluate three categories of MLLMs: (1) generalist models, including Molmo 7B (deitke2024molmo), LLaVA-1.6 7B (liu2024improved), Qwen2-VL 7B (wang2024qwen2), and Qwen2.5-VL 7B & 72B (bai2025qwen2); (2) math-specialized models, including Math-PUMA 7B (zhuang2024math), URSA 8B (luo2025ursa), and Math-LLaVA 13B (shi-etal-2024-math); and (3) proprietary models, GPT-4o (achiam2023gpt) and Gemini 2.0 Flash (google2024gemini2).
+Details of the experimental setup and prompts are provided in the Appendix.
+5.1 Toy: Solution Recognition
+The multimodal solution explanation tasks are designed to evaluate specific abilities rather than general problem-solving skills.
+To examine whether the model genuinely understands how to solve problems, we first perform a preliminary study on ME2 prior to the multimodal solution explanation task.
+Metrics.
+ME2 consists of both multiple-choice and short-answer problems. We report accuracy following the MathVista evaluation protocol (lu2023mathvista).
+Results.
+Table 2 shows the accuracy of problem-solving.
+The 7B generalist baseline struggles, while the 72B model performs second best.
+Somewhat unexpectedly, the math-specialized models perform worse than the generalist models, likely due to hindered instruction-following capabilities.
+Among the proprietary baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall.
+These results indicate that most MLLMs struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task.
 Problem-Solving (Acc)
@@
 Gemini 2.0 F - 0.481 0.291 0.436
-Table 2: Experimental results on the Solution Recognition toy task from ME2. The best scores are shown in bold, and the second-best scores are underlined.
-4 Experiments
-Models.
-We evaluate three categories of MLLMs: Open-source Generalist (OG), Open-source Mathematical specialist (OM), and Closed-source Generalist (CG).
-The OG models includes widely used models such as Molmo 7B (Deitke et al.
-2024), LLaVA-1.6 7B (Liu et al.
-2024), Qwen 2-VL (Wang et al.
-2024b), and Qwen 2.5-VL in both 7B and 72B versions (Bai et al.
-2025).
-The CG models comprise GPT-4o (Achiam et al.
-2023) and Gemini 2.0 Flash (Google DeepMind 2024).
-Finally, the OM models, specialized for mathematics, include Math-PUMA 7B (Zhuang et al.
-2024), URSA 8B (Luo et al.
-2025), and Math-LLaVA 13B (Shi et al.
-2024).
-Details of the experimental setup and prompts are provided in the Supplementary Materials.
-4.1 Toy: Solution Recognition
-Inspired by real-world tutoring settings where tutors typically know the solution path, the multimodal solution explanation tasks are designed under the assumption that the model has access to the correct solution, either in the form of a summary solution or a set of visual keypoints.
-To assess whether the model understands how to solve problems, we first conduct a preliminary study on ME2 before proceeding with the multimodal solution explanation task.
-Metrics.
-ME2 consists of both multiple-choice and short-answer problems. We report accuracy following the MathVista evaluation protocol (Lu et al. 2023).
-Results.
-Table 2 shows the accuracy of each model on the Solution Recognition toy task.
-The 7B OG baseline struggles, while the 72B model performs second best.
-Somewhat unexpectedly, the OM models perform worse than the OG models, likely due to aggressive fine-tuning that hinders instruction-following capabilities.
-Math-PUMA, despite being fine-tuned for mathematics on Qwen2-VL, also underperforms compared to the original models.
-Among the CG baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall.
-These results indicate that most open-source models struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task
+Table 2: Experimental results on the Solution Recognition toy task from ME2.
+Models are grouped into three categories: generalist models (top), math-specialized models (middle), and proprietary models (bottom).
+The best scores are in bold, and the second-best scores are underlined.
 5.2 Visual Keypoint Identification
 Metrics.
-We evaluate baselines’ VK identification performance using accuracy under a multiple-choice setting.
+We evaluate baseline performance using accuracy in a multiple-choice setting.
 Results.
-Table 3 summarizes performance on the VK Identification task.
-Among the OG baselines, the 7B models still face considerable difficulty, while the 72B model remains the second-best performer, similar to the Solution Recognition toy task.
-In contrast, the OM models perform at or below chance level (Acc = 0.20), even worse than on the toy task, suggesting a fundamental inability to detect relevant visual cues or a loss of instruction-following capability.
-Among the CG models, GPT-4o shows the most noticeable improvement despite its earlier underperformance on the toy task, while Gemini consistently achieves the highest performance.
-These findings indicate that although most open-source models still struggle despite having access to the solution, CG models and larger OG models can reasonably identify the visual keypoints.
+Table 3 summarizes performance on the visual keypoint identification task.
+The 7B generalist models struggle, while the 72B model remains the second-best performer.
+In contrast, math-specialized models perform near chance level (Acc = 0.20), indicating severe difficulty in identifying visual cues.
+Among proprietary models, Gemini achieves the highest performance.
+Overall, most models struggle to identify visual keypoints even with access to the solution, though proprietary ones perform relatively better.
+Since visual keypoint identification was evaluated under the assumption that models can already perform problem-solving, Table 4 reports success rates for each task and for both together.
+Only 23%, 10%, and 4% of proprietary, generalist, and math-specialized models succeed on both.
+This result highlights that real educational use remains challenging and that improving problem-solving ability is essential alongside visual keypoint identification.
 VK Identification (Acc)
@@
 Table 3: Experimental results for the Visual Keypoint Identification task on ME2, where models are evaluated on their ability to select the correct keypoints from multiple-choices.
+Model PS Only VKI Only PS $\cap$ VKI
+$\text{Qwen2.5-VL}^{\text{7B}}$ 18.9% 29.8% 10.5%
+Math-PUMA 19.7% 15.5% 4.6%
+Gemini 2.0 F 20.5% 34.5% 23.1%
+Table 4: Proportion (%) of cases where each model succeeds only on Problem-Solving (PS), only on Visual Keypoint Identification (VKI), or on both (PS $\cap$ VKI)
 Figure 5: Examples of reasoning processes and final predictions produced by Qwen2.5-VL 7B, Math-PUMA, and Gemini 2.0 Flash on the Visual Keypoint Identification task.
@@
 Each criterion is rated on a 5-point Likert scale.
-We report results from both human evaluators (Zheng et al.
-2023) and an LLM-based evaluator (Achiam et al.
-2023).
-In addition, we report text similarity metrics, including BLEU, ROUGE, METEOR, and BERTScore (Papineni et al.
-2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al.
-2019), with further details provided in the Supplementary Materials.
+We report results from both human evaluators (zheng2023judging) and an LLM-based evaluator using GPT-4o (achiam2023gpt).
+In addition, we report text similarity metrics, including BLEU, ROUGE, METEOR, and BERTScore (papineni2002bleu; lin2004rouge; banerjee2005meteor; zhang2019bertscore), with further details provided in the Appendix.
 Model Correctness Fidelity Referencing
@@
 Qwen2-VL 1.844 1.623 1.898
-[…]
-[…]
+$\text{Qwen2.5-VL}^{\text{7B}}$ 3.005 2.375 3.132
+$\text{Qwen2.5-VL}^{\text{72B}}$ 3.397 3.048 3.533
 Math-PUMA 1.770 1.609 1.581
@@
 Table 5: LLM-based evaluation results for the Keypoint-based Explanation Generation task on ME2, rated on a 1-5 Likert scale across three criteria: (1) Correctness, assessing logical validity; (2) Fidelity, measuring alignment with the intent of the reference explanation; and (3) Referencing, evaluating the appropriate use of key visual elements.
+Results.
+Table 5 presents the results of the explanation generation task.
+While most models achieve reasonable Correctness, many fail to follow the intended reasoning path (Fidelity) or reference the given keypoints (Referencing).
+Generalist models struggle overall, though the Qwen2.5-VL series shows size-dependent improvement.
+Math-specialized models still fail to produce coherent or instruction-following explanations.
+In contrast, proprietary models achieve the highest scores, demonstrating stronger abilities in generating well-grounded explanations.
+As shown in Table 6, human evaluation exhibits strong correlations with LLM judgments, as indicated by the Spearman coefficients (zar2005spearman) (0.770 for Correctness, 0.783 for Fidelity, 0.788 for Referencing; all p<<0.05).
+These results show that although most open-source models still struggle, proprietary models and more recent generalist models can generate appropriate explanations.
+6 Analyses
+6.1 Qualitative Analysis
+To analyze how the three categories of models differ in their outputs, we examined the results from representative models in each category: Qwen2.5-VL 7B (generalist), Math-PUMA (specialist), and Gemini 2.0 Flash (proprietary).
 Model Correctness Fidelity Referencing
-[…]
+$\text{Qwen2.5-VL}^{\text{7B}}$ 2.610 2.541 2.610
 Math-PUMA 1.041 1.041 1.037
@@
 Table 6: Human evaluation results for the Keypoint-based Explanation Generation task, rated on a 1–5 Likert scale.
-Results.
-Table 4 presents the results of the explanation generation task.
-Although all models perform relatively well in terms of Correctness by aligning their explanations with the problem, many fail to follow the intended reasoning path (Fidelity) or to reference the given visual keypoints (Referencing).
-OG models generally struggle, whereas the Qwen2.5-VL series performs relatively well, showing improved performance with increasing model size.
-OM models still fail to generate coherent explanations or follow instructions.
-In contrast, CG models achieved the highest scores in this explanation task, indicating their strong capabilities in generating well-grounded explanations.
-As shown in Table 5, human evaluation results show a similar pattern, with strong correlations between human and LLM judgments as indicated by the Spearman coefficients (Zar 2005), which are 0.770 for Correctness, 0.783 for Fidelity, and 0.788 for Referencing, all statistically significant (p$<$0.05).
-These results suggest that while most open-source models still struggle, CG models and recent OG models like the Qwen2.5-VL series are capable of generating appropriate explanations.
-5 Analyses
-5.1 Qualitative Analysis
-To analyze how the three categories of models differ in their outputs, we examined the results from representative models in each category: Qwen2.5-VL 7B (OG), Math-PUMA (OM), and Gemini 2.0 Flash (CG).
+Figure 6: Error analysis for Visual Keypoint Identification: Correct choice rates, (1) cases where the correct element was chosen but referenced incorrectly, (2) choices containing more keypoints than required, and (3) choices containing fewer keypoints than needed.
 Visual Keypoint Identification
-Figure 5 shows VK identification examples from three model categories.
-The Qwen2.5-VL tries to reason about which option provides the most informative visual keypoints for the task, yet its final selection is wrong.
-The Math-PUMA shows neither coherent reasoning nor a correct answer.
-In contrast, the Gemini 2.0 Flash quickly grasps the core of the instruction, analyzes candidate visual keypoints, and selects the option that best supports understanding the problem.
+Figure 5 shows visual keypoint identification examples from three model categories.
+Qwen2.5-VL attempts to reason about the most informative keypoints but still selects the wrong option.
+Math-PUMA shows neither coherent reasoning nor a correct answer.
+In contrast, Gemini 2.0 Flash correctly interprets the instruction, analyzes the candidates, and chooses the most appropriate keypoints.
+Overall, similar patterns were consistently observed across examples.
+To gain a finer-grained understanding of the models, we analyze failure cases on a 10% subset of the dataset, as shown in Figure 6.
+We categorize errors into three types: (1) electing the correct elements but providing incorrect descriptions, (2) selecting options that contain extra elements, and (3) selecting options with missing required elements.
+The three models exhibit similar rates for incorrect descriptions and extra elements but differ substantially in missing elements: Math-PUMA shows the highest rate (38%), followed by Qwen2.5-VL (20%).
+In contrast, Gemini 2.0 Flash achieves the highest accuracy, while Qwen2.5-VL performs moderately, and Math-PUMA remains the least reliable among the three.
 Figure 7: Examples of explanations from three models on the Keypoint-based Explanation Generation task, along with their evaluated scores in Correctness (C), Fidelity (F), and Referencing (R).
@@
 Figure 7 presents explanation examples from three model categories.
-In this example, Qwen2.5-VL scored 2 for Correctness, 2 for Fidelity, and 3 for Referencing.
-Although the model is given keypoints, its approach is not aligned with the reference solution.
-Its explanation references visual cues such as angles, lengths, symbols, and triangles, which indicates partial alignment with the solution image.
-However, it ultimately misinterprets the problem and produces incorrect reasoning.
+In this example, Qwen2.5-VL scored 2 in Correctness, 2 in Fidelity, and 3 in Referencing.
+Despite being provided with keypoints, its reasoning diverges from the reference and only partially aligns with the solution image, ultimately leading to an incorrect interpretation.
 Math-PUMA received a score of 1 across all metrics, as it only returned the final answer without any supporting explanation.
-In contrast, Gemini scored 5 in all metrics.
-It generated an explanation as if the keypoints were embedded in the solution image, following the same reasoning structure as the reference and correctly referring to the same key visual elements throughout the solution.
-Model PS Only VKI Only PS $\cap$ VKI
-[…]
-[…]
-Math-PUMA 19.7% 15.5% 4.6%
-GPT-4o 14.6% 35.9% 11.3%
-Gemini 2.0 F 20.5% 34.5% 23.1%
-Table 6: Proportion (%) of cases where each model succeeds only on Problem-Solving (PS), only on Visual Keypoint Identification (VKI), or on both (PS $\cap$ VKI)
-5.2 Problem Solving and VK Identification
-Table 6 reports the proportion of cases where models succeed only on Problem-Solving (PS), only on Visual Keypoint Identification (VKI), or on both.
-Overall, generalist models tend to perform better on VKI than on PS alone, suggesting that most models are more capable of recognizing keypoints assuming the reasoning path is known, rather than solving the problems end-to-end.
-Notably, Qwen 72B and Gemini achieve over 20% accuracy in cases involving both PS and VKI, demonstrating relatively strong performance.
-However, since current models still struggle with problem-solving, improving multimodal mathematical reasoning is essential for advancing performance on both PS and VKI.
+In contrast, Gemini scored 5 across all metrics, generating an explanation that mirrored the reference reasoning and consistently referred to the correct key visual elements.
+Similar to the experimental results in Figure 6, proprietary models generate coherent explanations, whereas open-source models struggle, with math-specialized ones showing almost no reasoning capability.
 Figure 8: Attention maps from various layers of open-source MLLMs on the Visual Keypoint Identification task.
@@
 6.2 Do MLLMs Attend to Visual Keypoints?
-To analyze whether current open-source models can accurately identify visual keypoints in images, we examined the attention maps of OG and OM models on the multimodal solution explanation tasks.
-Figure 7 illustrates the attention patterns of the OG and OM models.
-Both models show a tendency to attend to both global and local regions of the input image.
-However, despite the explicit inclusion of visual keypoints in the options and summary solution of the Visual Keypoint Identification task, neither model effectively focuses on the relevant visual regions.
+To analyze how current open-source models fail to visually recognize keypoints, we examined the attention maps of both generalist and math-specialized models on the Visual Keypoint Identification task.
+Figure 8 illustrates the attention patterns, showing that both models tend to attend well to both global and local regions of the input image.
+However, despite the explicit inclusion of visual keypoints in the options and summary solution, neither model effectively focuses on the relevant visual regions.
 A similar pattern is observed in the Keypoint-based Explanation Generation task.
-These findings suggest that current open-source models still struggle to visually attend to the keypoints, even when they are essential for understanding or explaining the problem.
-6 Related Works
-Language Models for Education.
-Recent advances in Large Language Models (LLMs) (Brown et al.
-2020) have sparked significant interest in educational applications, particularly in personalized problem recommendation (Park et al.
-2024), automated tutoring (Chevalier et al.
-2024), and the provision of tailored feedback and customized curricula (Hu et al.
-2024; Macina et al.
-2023; Feng, Wang, and Sun 2023).
-For effective education, research suggests that combining textual and visual information enhances comprehension and memory more effectively than using text alone (Arcavi 2003; Stylianou 2010; Lee, Park, and Park 2024).
-As Clark and Paivio (1991) explains, dual representations supply multiple retrieval cues and cultivate richer mental models.
-Building on this insight, we introduce the multimodal solution explanation task to enable LLMs to offer learners more comprehensive learning opportunities.
-This task enables the model to pinpoint the visual keypoints essential for students’ comprehension and to generate explanations grounded in those keypoints, thereby delivering more comprehensive educational support and ultimately improving the overall quality of their learning experience.
-Multimodal Mathematical Benchmarks.
-Current LLMs show strong performance on mathematical problems, making them valuable tools for students (Zhuang et al.
-2024; Luo et al.
-2025).
-To evaluate these models, traditional mathematical benchmarks (Cobbe et al.
-2021; Hendrycks et al.
-2021) have been crucial in assessing reasoning capabilities.
-With the growing multimodal capabilities of LLMs, benchmarks such as MathVista (Lu et al.
-2023), Math-Vision (Wang et al.
-2024a), and MathVerse (Zhang et al.
-2024) have extended this evaluation to image-based math problems.
-Recent efforts like OlympiadBench (He et al.
-2024) and MM-MATH (Sun et al.
-2024) further assess models not just on final answers but also on their reasoning processes.
-However, most existing benchmarks focus solely on problem-solving, overlooking educational objectives.
-To address this gap, we introduce ME2, a new benchmark designed to evaluate a model’s multimodal solution explanation capability.
-Together, ME2 and the multimodal solution explanation task are expected to advance educational research and foster the development of more effective teaching models.
-7 Disscusion and Conclusion
-We introduce the multimodal solution explanation task, which requires core mathematics-teaching capabilities for generating intuitive and insightful explanations.
-To benchmark these capabilities, we propose ME2 benchmark, which consists of two subtasks: (1) Visual Keypoint Identification, where models must select visual keypoints that are essential for understanding and explaining the problem, and (2) Keypoint-based Explanation Generation, which requires models to produce explanations grounded in those visual keypoints that support intuitive reasoning.
-Our experimental results show that current MLLMs struggle even with the problem-solving aspects of the ME2 tasks.
-Furthermore, even under the assumption that the model already knows the solution path, similar to a human tutor, we found that most open-source models still fail to accurately identify visual keypoints, with the exception of closed-source models and recent open-source models.
-Surprisingly, even math-specialized models perform worse in this aspect.
-The performance gap becomes even more pronounced in the explanation generation task.
-To perform well on the multimodal solution explanation task, future open-source models may need the following advancements:
-  1. 1.
-    Enhanced visual grounding mechanisms that link visual elements to multimodal mathematical reasoning.
-  2. 2.
-    Advanced solution-path tracking capabilities for multimodal explanations.
-We believe this work will inspire future research along two key paths: (1) enhancing models’ mathematical reasoning capabilities through improved visual grounding, and (2) leveraging visual grounding in educational contexts to develop AI systems that can serve as real-world mentors, ultimately providing clear, multimodal, and student-friendly explanations.
-The code and data will be released publicly.
+While current MLLMs are capable of attending to general visual content (zhang2025mllms), our analysis suggests that they still lack robust mathematical visual grounding ability.
+Improving visual grounding in mathematical contexts will likely be essential for future models to perform well on the multimodal solution explanation task.
+7 Conclusion
+We introduce the multimodal solution explanation task and the ME2 benchmark, which assess multimodal mathematics‐teaching capabilities through two complementary subtasks: identifying essential visual keypoints and generating explanations grounded in those keypoints.
+Our experimental results demonstrate that current models struggle with both problem solving and visual keypoint identification, and that this performance gap becomes more pronounced for open-source models in the explanation generation task.
+Through our analysis, we find that this limitation stems from the lack of math-specific visual grounding and robust visually grounded reasoning.
+Enhancing these two abilities will be essential for applying MLLMs effectively in educational settings.
+8 Ethics Statement
+In this paper, we introduce the ME2 benchmark, curated from an in-house mathematics education platform [^1] https://mathpresso.com/en.
+All materials were reviewed to ensure full compliance with copyright and data usage regulations.
+Data annotation was conducted by bilingual annotators holding undergraduate degrees in relevant fields, ensuring adequate mathematical and linguistic proficiency.
+9 Acknowledgments
+This work was partly supported by an Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No.
+RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2025-02263598, Development of Self-Evolving Embodied AGI Platform Technology through Real-World Experience), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00354218, RS-2024-00353125).
+We express special thanks to KAIT GPU project.
+The ICT at Seoul National University provides research facilities for this study.
+Explain with Visual Keypoints Like a Real Mentor!
 A Benchmark for Multimodal Solution Explanation
-Supplementary Material
-Appendix A Overview
-This supplementary material includes the following sections: (1) Detailed experimental settings used in our benchmark, (2) Details of our benchmark, and (3) Prompts used for model response generation
-Appendix B Experiment Details
-In this study, experiments were conducted using the LMMs-eval repository (Li et al. 2024). This repository provides a comprehensive framework for evaluating multi-modal models across various tasks.
-B.1 Evaluated Models
-In all LLM-based evaluations, we used the gpt-4o-2024-08-06 endpoint.
+Technical Appendix
+Appendix A Experiment Details
+In this study, experiments were conducted using the LMMs-eval repository  [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval.
+This repository provides a comprehensive framework for evaluating multi-modal models across various tasks.
+A.1 Computational Resources
+For closed-source models, we used the OpenAI API and Gemini Developer API to infer the output of GPT-4o and Gemini 2.0.
+For open-source models such as Math-PUMA, URSA, MathLLaVA, LLaVA, Qwen-2-VL, Qwen-2.5-VL, and Molmo, inference was performed using a NVIDIA A6000 48GB GPU.
+While the exact inference speed varies depending on the task, model, and lengths of prompts and responses, a query takes about 50 seconds to be answered.
+A.2 Evaluated Models
 When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^3] https://huggingface.co/models.
@@
 The selection of models spans a range of architectures and performance levels, providing insights into current advancements in multi-modal learning.
-B.2 Computational Resources
-For closed-source models, we used the OpenAI API and Gemini Developer API to infer the output of GPT-4o and Gemini 2.0.
-For open-source models such as Math-PUMA, URSA, MathLLaVA, LLaVA, QWEN-2-VL, QWEN-2.5-VL, and Molmo, inference was performed using a NVIDIA A6000 48GB GPU.
-While the exact inference speed varies depending on the task, model, and lengths of prompts and responses, a query takes about 50 seconds to be answered.
-B.3 Human Evaluation Setup
+A.3 LLM Evaluation
+In all LLM-based evaluations, we used the gpt-4o-2024-08-06 endpoint.
+For the Keypoint-based Explanation Generation task, we compared the rankings obtained when Math-PUMA and GPT outputs were evaluated separately by Gemini and GPT.
+The resulting Kendall’s $\tau$ values were 0.90 for Correctness, 0.84 for Fidelity, and 0.92 for Referencing.
+Although evaluation bias is often a concern when an LLM assesses models from the same family, the high agreement between the GPT-judge and Gemini-judge indicates that no substantial bias is present.
+A.4 Human Evaluation
 Three evaluators, all holding a bachelor’s or master’s degree in engineering, are assessing AI model outputs for 80 problems.
-Specifically, they are evaluating the results produced by the Math-PUMA, Qwen2.5-VL, and Gemini 2.0 Flash models, which represent the OM, OG, and CG categories.
+Specifically, they are evaluating the results produced by the Math-PUMA, Qwen2.5-VL, and Gemini 2.0 Flash models, which represent the math-specialized, generalist, and proprietary models categories.
 Each criterion is being rated on a five-point Likert scale.
-B.4 Evaluation on automatic metrics
-We report additional evaluation results for Keypoint-based Explanation Generation task using automatic metrics, including BLEU, ROUGE, METEOR, and BERTScore (Papineni et al.
-2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al.
-2019).
+LLM–human agreement shows strong correlations, as indicated by the Spearman coefficients (0.770 for Correctness, 0.783 for Fidelity, and 0.788 for Referencing; all $p<0.05$).
+Human–Human agreement, measured by Krippendorff’s $\alpha$, reached 0.696 for Correctness, 0.571 for Fidelity, and 0.612 for Referencing.
+A.5 Automatic Metrics Evaluation
+We report additional evaluation results for Keypoint-based Explanation Generation task using automatic metrics, including BLEU, ROUGE, METEOR, and BERTScore.
 The detaileed scores can be found in Table A.
-Appendix C Benchmark Details
+Appendix B Benchmark Details
 Our ME2 benchmark consists of a total of 17 chapters and 33 sections as shown in Table B.
-Appendix D Prompts
-This section compiles all the prompts used in our experiments.
-The prompts shown in Figure A, Figure B, and Figure C are used to generate model outputs for the Solution Recognition toy task, Visual Keypoint Identification, and Keypoint-based Explanation Generation tasks, respectively.
-For the Keypoint-based Explanation Generation task, model responses are evaluated using the prompts in Figure D.
-Appendix E Ethical Considerations
-In this paper, we introduce the ME2 benchmark, curated from an in-house mathematics education platform.
-All materials were reviewed to ensure compliance with copyright regulations.
-Data annotation was performed by bilingual annotators with undergraduate degrees, ensuring sufficient mathematical problem-solving skills.
+During dataset validation, we also reviewed the options related to visual keypoint identification and confirmed their consistency and reliability.
+Appendix C Ablation on Solution Summary Anchoring
+Since a problem may have multiple valid solutions and the corresponding visual keypoints can vary, we provide the model with a solution summary ($T_{s}^{tldr}$) that anchors a single explanatory direction.
+To examine the effect of this anchoring, we analyze the qualitative differences when the summary is not provided in Figure A.
+As shown, without anchoring, the model’s explanations often drift toward alternative reasoning paths or focus on irrelevant keypoints.
+Figure A: Comparison of explanations generated with and without the solution summary ($T_{s}^{tldr}$).
+Without anchoring, models often drift to alternative reasoning paths or irrelevant keypoints.
+(a) Without $T_{s}^{tldr}$; (b) With $T_{s}^{tldr}$.
 Model Params BLEU-2 BLEU-4 ROUGE-L METEOR BERTScore
-OG Molmo 7B 0.158 0.067 0.187 0.310 0.842
+Molmo 7B 0.158 0.067 0.187 0.310 0.842
 LLaVA-1.6 7B 0.130 0.059 0.176 0.287 0.835
@@
 Qwen2.5-VL 72B 0.097 0.043 0.190 0.316 0.821
-OM Math-PUMA 7B 0.006 0.002 0.119 0.058 0.818
+Math-PUMA 7B 0.006 0.002 0.119 0.058 0.818
 URSA 8B 0.020 0.006 0.079 0.075 0.735
 Math-LLaVA 13B 0.112 0.057 0.147 0.252 0.817
-CG Gemini 2.0 Flash - 0.149 0.083 0.210 0.367 0.842
+Gemini 2.0 Flash - 0.149 0.083 0.210 0.367 0.842
 GPT-4o - 0.095 0.045 0.161 0.301 0.815
 Table A: Experimental results of automated evaluation metrics for the Keypoint-based Explanation Generation task on ME2.
+Appendix D Prompts
+This section compiles all the prompts used in our experiments.
+The prompts shown in Figure B, Figure C, and Figure D are used to generate model outputs for the Solution Recognition toy task, Visual Keypoint Identification, and Keypoint-based Explanation Generation tasks, respectively.
+For the Keypoint-based Explanation Generation task, model responses are evaluated using the prompts in Figure E.
 Chapter Title Section Title
@@
 Your final response must END with the format.
-Figure A: Prompt used for the Visual Keypoint Identification task in the ME2 benchmark. Prompt inputs are boldfaced.
+Figure B: Prompt used for the Visual Keypoint Identification task in the ME2 benchmark. Prompt inputs are boldfaced.
 Q: As shown in the figure, there are 5 points: A, B, C, D, and E.
@@
 solution_text:
-Figure B: Prompt used for the Keypoint-based Explanation Generation task in the ME2 benchmark.
+Figure C: Prompt used for the Keypoint-based Explanation Generation task in the ME2 benchmark.
 It is designed to generate educationally effective explanations for the given math problem.
@@
 Find the value of a + b.
-Again, only return the final answer.
-Any additinal text will be considered incorrect.
-Figure C: Prompt used for the Solution Recognition toy task in the ME2 benchmark. It is designed to generate an answer to the given math problem. Prompt inputs are boldfaced.
+Again, only return the final answer. Any additinal text will be considered incorrect.
+Figure D: Prompt used for the Solution Recognition toy task in the ME2 benchmark. It is designed to generate an answer to the given math problem. Prompt inputs are boldfaced.
 You are evaluating the quality of an AI-generated explanation for a math problem involving geometry or graph-based reasoning.
@@
 - Rate on a Likert scale: **1, 2, 3, 4, or 5**
-4. Mathematical Accuracy
-- Does the model produce the same final answer as the reference?
-- If the model’s answer is correct, give a score of **1**.
-- If the model’s answer is incorrect, give a score of **0**.
 ### Output Format
 Important: Report your rating using the exact format below:
-Rating: [[x, y, z, w]]
-— where ‘x‘ is your score for correctness, ‘y‘ for reference alignment, ‘z‘ for use of visual elements, and ‘w‘ for mathematical accuracy.
-Figure D: GPT evaluation prompt used to assess model outputs for the Keypoint-based Explanation Generation task in ME2. Prompt inputs are boldfaced.
+Rating: [[x, y, z]]
+— where ‘x‘ is your score for correctness, ‘y‘ for reference alignment, and ‘z‘ for use of visual elements.
+Figure E: GPT evaluation prompt used to assess model outputs for the Keypoint-based Explanation Generation task in ME2. Prompt inputs are boldfaced.

### v5

@@
 Archivehttps://me2-benchmark.github.io
-Figure 1: A student solving a math problem often benefits from visual cues, such as lines, symbols, or highlights, that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions.
+Figure 1: A student solving a math problem often benefits from visual cues—such as lines, symbols, or highlights—that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions.
 To serve as effective educational assistants, machines must go beyond answer generation and emulate human-like explanation strategies by explicitly incorporating and referencing visual elements.
@@
    - Short-answer questions 96
-Average number of VKVK 3.8
+Average number of VKVK 3.73
 Maximum words in $T_{p}$ 211
@@
 Maximum words in $vk_{n}$ 45
-Maximum words in $T_{s}^{tldr}$ 123
+Maximum words in $T_{s}^{tldr}$ 198
 Average words in $T_{p}$ 53.1
-Average words in $T_{s}$ 102.4
+Average words in $T_{s}$ 101.4
 Average words in $vk_{n}$ 12.2
-Maximum words in $T_{s}^{tldr}$ 35.7
+Maximum words in $T_{s}^{tldr}$ 35.8
 Table 1: Statistics of the ME2 benchmark, including problem subjects, types, and instance word counts.
@@
 The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem.
-We compute text similarity using Qwen3 Embedding (qwen3embedding).
+Text similarity was computed using Qwen3 Embedding (qwen3embedding), and to ensure reliability, human annotators carefully reviewed and corrected any options exhibiting logical inconsistencies.
 Keypoint-based Explanation Generation.
@@
 The multimodal solution explanation tasks are designed to evaluate specific abilities rather than general problem-solving skills.
-To examine whether the model genuinely understands how to solve problems, we first perform a preliminary study on ME2 prior to the multimodal solution explanation task.
+To examine whether the model genuinely understands how to solve problems, we first perform a preliminary study on ME2 prior to the multimodal explanation task.
 Metrics.
@@
 Among the proprietary baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall.
-These results indicate that most MLLMs struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task.
+These results indicate that most MLLMs struggle to recognize the correct solution on ME2, even before performing the multimodal explanation task.
 Problem-Solving (Acc)
@@
 Table 6: Human evaluation results for the Keypoint-based Explanation Generation task, rated on a 1–5 Likert scale.
-Figure 6: Error analysis for Visual Keypoint Identification: Correct choice rates, (1) cases where the correct element was chosen but referenced incorrectly, (2) choices containing more keypoints than required, and (3) choices containing fewer keypoints than needed.
-Visual Keypoint Identification
+Figure 6: Error analysis for Visual Keypoint Identification: (1) cases where the correct element was chosen but referenced incorrectly, (2) choices containing more keypoints than required, and (3) choices containing fewer keypoints than needed.
+Visual Keypoint Identification.
 Figure 5 shows visual keypoint identification examples from three model categories.
@@
 Overall, similar patterns were consistently observed across examples.
-To gain a finer-grained understanding of the models, we analyze failure cases on a 10% subset of the dataset, as shown in Figure 6.
-We categorize errors into three types: (1) electing the correct elements but providing incorrect descriptions, (2) selecting options that contain extra elements, and (3) selecting options with missing required elements.
+To gain a finer-grained understanding of the models, we analyze their output behaviors on a 10% subset of the dataset, as shown in Figure 6.
+We categorize all incorrect outputs into three types: (1) electing the correct elements but providing incorrect descriptions, (2) selecting options that contain extra elements, and (3) selecting options with missing required elements.
 The three models exhibit similar rates for incorrect descriptions and extra elements but differ substantially in missing elements: Math-PUMA shows the highest rate (38%), followed by Qwen2.5-VL (20%).
@@
 Qwen2.5-VL starts with valid yet unaligned explanations, which soon become incorrect; Math-PUMA generates no explanation; Gemini 2.0 Flash generates solution-aligned explanations.
-Keypoint-based Explanation Generation
+Keypoint-based Explanation Generation.
 Figure 7 presents explanation examples from three model categories.
@@
 We introduce the multimodal solution explanation task and the ME2 benchmark, which assess multimodal mathematics‐teaching capabilities through two complementary subtasks: identifying essential visual keypoints and generating explanations grounded in those keypoints.
-Our experimental results demonstrate that current models struggle with both problem solving and visual keypoint identification, and that this performance gap becomes more pronounced for open-source models in the explanation generation task.
+Our experimental results demonstrate that current models struggle with both problem-solving and visual keypoint identification, and that this performance gap becomes more pronounced for open-source models in the explanation generation task.
 Through our analysis, we find that this limitation stems from the lack of math-specific visual grounding and robust visually grounded reasoning.
 Enhancing these two abilities will be essential for applying MLLMs effectively in educational settings.
-8 Ethics Statement
+Ethics Statement
 In this paper, we introduce the ME2 benchmark, curated from an in-house mathematics education platform [^1] https://mathpresso.com/en.
@@
 Data annotation was conducted by bilingual annotators holding undergraduate degrees in relevant fields, ensuring adequate mathematical and linguistic proficiency.
-9 Acknowledgments
+Acknowledgments
 This work was partly supported by an Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No.